OpenStudy (kitkat1):

Mr. James has cylindrical beakers that measure 4 inches in diameter and are 9 inches high. What is the volume contained within the beakers? Use 3.14 for pi. Round your answer to the nearest hundredth. 226.08 cubic inches 113.04 cubic inches 56.52 cubic inches 100.34 cubic inches

OpenStudy (kitkat1):

Please someone help me really need help will give out medals..

OpenStudy (anonymous):

\[V = \pi r^2 h\] r = diameter / 2 h = 9 pi = 3.14

OpenStudy (kitkat1):

So V=3.14) r^2) 9

OpenStudy (kitkat1):

and what do I have to multiply here there is no radius?

OpenStudy (kitkat1):

So if mulitplying without the radius its 56.52 are u sure about that?

OpenStudy (anonymous):

diameter = 2* radius

OpenStudy (kitkat1):

so what is the answer it may be 56/52 or 113.04

OpenStudy (kitkat1):

I think

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

113.01

OpenStudy (kitkat1):

so whats the answer?

OpenStudy (kitkat1):

okay 113.01

OpenStudy (anonymous):

\[V = 3.14 *(2)^2 * 9 = 3.14 *4* 9 = 113.0\]

OpenStudy (kitkat1):

A pillar candle is shaped like a cylinder with a radius of 1.5 inches. The candle holds 123 cubic inches of wax. How tall is the candle? Use 3.14 for pi. Round your answer to the nearest hundredth. (4 points) 12.50 inches 17.41 inches 18.76 inches 15.76 inches

OpenStudy (anonymous):

Volume = 123 cubic in radius = 1.5 in find h using \[V = \pi r^2 h\]

OpenStudy (kitkat1):

v=3.14) 1.5)^2(h)

OpenStudy (kitkat1):

How can i find the Height?

OpenStudy (anonymous):

123 = (3.14)* (1.3)^2 * h h = 123/ (3.14)*(1.5)^2

OpenStudy (kitkat1):

so it should be 17.41?

OpenStudy (anonymous):

yes.

OpenStudy (kitkat1):

can u help me with one more please An average clementine is shaped like a sphere and has a diameter of 3.5 inches. What is the volume of the fruit? Use 3.14 for pi. Round your answer to the nearest hundredth. (4 points) 23.78 cubic inches 22.44 cubic inches 21.34 cubic inches 20.45 cubic inches

OpenStudy (kitkat1):

nevermind i found that the answer is 22 44

OpenStudy (kitkat1):

Thanks

OpenStudy (anonymous):

anytime :)